The Bruin's Den Outdoor Gear is considering a new 7-year project to produce a new tent line. The equipment necessary would cost $1.67 million and be depreciated using straight-line depreciation to a book value of zero. At the end of the project, the equipment can be sold for 10 percent of its initial cost. The company believes that it can sell 27,000 tents per year at a price of $71 and variable costs of $31 per tent. The fixed costs will be $465,000 per year. The project will require an initial investment in net working capital of $221,000 that will be recovered at the end of the project. The required rate of return is 11.4 percent and the tax rate is 40 percent. What is the NPV?
Answer #1

Depreciation = Cost of the Project / Useful life years = $1,670,000 / 7 = $238,571.43

Cash flow at year 0 = Cost of Project + Investment in NWC = $1,670,000 + $221,000 = $1,891,000

Annual Cash Flow(Year 1 - 7) = [{(P - VC) * Qty. Sold} - Fixed Cost] * [1 - t] + [Depreciation * t]

= [{($71 - $31) * 27,000} - $465,000] * [1 - 0.40] + [$238,571.43 * 0.40]

= $369,000 + $95,428.57 = $464,428.57

Additional Cash Flow at year 7 = After-tax Salvage Value + Recovery of NWC

= [($1,670,000 * 0.10) * (1 - 0.40)] + $221,000 = $100,200 + $221,000 = $321,200

NPV = PV of Cash Inflows - PV of Cash Outflows

= [$464,428.57 * {1 - (1 + 0.114)-7} / 0.114] + [$321,200 / (1 + 0.114)7] - $1,891,000

= [$464,428.57 * 4.6519] + [$321,200 / 2.1291] - $1,891,000

= $2,160,482.09 + $150,861.77 - $1,891,000 = $420,343.86